# On the accuracy of posterior recovery with neural network emulators

### v1

@@ 3.5 Likelihood functions
where $N_{d}$ is the number of data points, $D$ is the mock data containing the fiducial ARES signal and Gaussian random noise with standard deviation $\sigma$, $\mathcal{M}$ is the model of the 21-cm signal from ARES or the emulated version from globalemu  [^2] The log-evidences found for the fits in this paper differ to those found in DJ23 by around 2000 log units. We do not report them here but they can be found with the chains and the code on github. We believe the difference is because we included the normalisation term on our log-likelihood, $-\frac{N_{d}}{2}\log(2\pi\sigma^{2})$, and DJ23 did not. For $N_{d}=490$ and $\sigma=25$ mK then this term is equal to approximately $-2027$. . We are assuming that the errors are not correlated across frequency, which is a common assumption in the field and was made in DJ23.
Usually, the magnitude of the noise in ones’ data is not known. Theoretical assumptions can be made about the form and magnitude of the noise using eq. 36, however, we often fit $\sigma$ as a free parameter in the analysis. Recent works (Gessey-Jones et al., 2024; Pochinda et al., 2024) have included an additional emulator error term in their likelihood functions, such that $\sigma^{2}=\sigma_{\mathrm{instrument}}^{2}+\sigma_{\mathrm{emulator}}^{2}$. The emulator error is often estimated from the test data sets and fixed in the analysis, whereas the instrument error is fitted for. For a sky-averaged 21-cm experiment, the instrument noise is expected to be around 25 mK (e.g. Bowman et al., 2018; de Lera Acedo et al., 2022) and emulator errors are typically around 1 mK and therefore $\sigma$ is generally dominated by the contribution from the instrument.
@@ 3.6 KL divergence
We use normalising flows (NFs) implemented with margarine (Bevins et al., 2022a, 2023) to evaluate the log-probabilities $\log P$ and $\log P_{\epsilon}$ for samples $\theta\sim P(\theta|D,\mathcal{M})$. NFs are a class of generative density estimation tools that use neural networks to parameterize invertible transformations between a known distribution such as a multivariate standard normal and a more complex target distribution like $P$ and $P_{\epsilon}$. Once trained, the flows can be used to generate samples on the distributions, i.e. $\theta\sim P(\theta|D,\mathcal{M})$, and via the change of variables formula calculate log-probabilities. Errors on the KL estimates are calculated using the method detailed in Bevins et al. (2023).
For each normalising flow used in this work, we have a learning rate of $10^{-3}$ and five neural networks chained together, each with one hidden layer of 250 nodes. We train for a maximum of 1000 epochs and early stop if the validation loss has not decreased after 20 epochs, rolling back to the optimum model. For more details on the particular implementation of normalising flows used in this work see Bevins et al. (2022a, 2023).
@@ 3.7 Results
Using the trained emulator, the likelihood function with a fixed noise and margarine to calculate the KL divergence, as discussed in the previous sections, we are able to compare the posteriors recovered when using ARES and globalemu to model the signal. We use the nested sampling implementation Polychord (Handley et al., 2015b, a) to sample the likelihood function. In all the fits presented, we use the default Polychord settings [^3] See https://github.com/htjb/validating_posteriors. and we show in appendix C that this is sufficient to recover consistent results. As in DJ23, we test recovery when fitting a fiducial 21-cm signal with Gaussian distributed noise that has a standard deviation of 5, 25, 50 and 250 mK. We show the posteriors for 25 mK noise in fig. 2 and for 5, 50 and 250 mK in appendix A.
@@
From fig. 1, we might expect the recovery of the 1D posteriors on $f_{\mathrm{esc}}$, $C_{X}$ and $\log N_{\mathrm{HI}}$ to be worse than other parameters as the emulator error is worse at lower redshifts. However, we find that the posteriors on these parameters are very consistent, suggesting that the performance of the emulator even in this part of the parameter space is good. We report the ‘emulator bias’ for each parameter and each level of noise in appendix B. We find that the emulation bias is consistently below 1 for all the parameters and all the noise levels. At most, the emulation bias is 0.69 for $\log T_{\mathrm{min}}$ when $\sigma=5$ mK and we find that the average emulation bias decreases with increasing noise from 0.33 to 0.04 at 250 mK.
@@
Using eq. 32 we can estimate an approximate upper limit on the $\mathcal{D}_{\mathrm{KL}}$ between the two posteriors for a given level of noise in the data and a given emulator error. In table 3 we report these limits using the mean and 95th percentile RMSE values for the emulator for each level of noise considered in this paper. Since the 95th percentile RMSE is larger than the mean value and the limit is $\propto\mathrm{RMSE}^{2}$ it gives a more conservative limit on the accuracy of the emulated posterior. We also estimate the $\mathcal{D}_{\mathrm{KL}}$ between the posteriors using the method described in section 3.6 and report these values in table 3. We see that for all the different noise levels the $\mathcal{D}_{\mathrm{KL}}$ are approximately consistent, within uncertainty, with 0 and consistent with the upper bounds from eq. 32.
@@
In fig. 3 we show how the upper limit on $\mathcal{D}_{\mathrm{KL}}$ changes with the magnitude of the noise in the data and the RMSE error on the emulator as a series of blue contours. We also show contours for the mean and 95th percentile RMSE values for the emulator used in this paper as red and green dashed lines. The vertical dotted lines mark the 5, 25, 50 and 250 mK noise values and the intersection between these lines and the dashed red and green lines give the predicted upper limits on the $\mathcal{D}_{\mathrm{KL}}$ reported in table 3. The purple scatter points give the $\mathcal{D}_{\mathrm{KL}}$ values, estimated with margarine, for each pair of posteriors. Although not perfect, it is clear that the limit in eq. 32 provides an approximate estimate for the maximum value of the $\mathcal{D}_{\mathrm{KL}}$ we might expect for a given noise and RMSE, and hence a good guide on the accuracy required of emulators for inference. For each pair of posteriors, the estimated $\mathcal{D}_{\mathrm{KL}}$ is approximately consistent with zero.
@@
Figure 2: The posteriors recovered when fitting the fiducial ARES signal directly with ARES in blue and globalemu in orange for $\sigma=25$ mK. The lower half of the triangle plot shows a kernel density estimation (KDE) of the 2D marginalised posteriors, the diagonal shows the 1D KDEs and the upper half shows the samples. We show the fiducial parameter values as red dashed lines for reference, but stress that we are more interested in the similarity between the posteriors in this work. While there are some small difference between the posteriors, they are visually similar. However, given the mean RMSE for the emulator and the limit outlined in eq. 32 this is not so surprising with the maximum predicted $\mathcal{D}_{\mathrm{KL}}$ between the emulated and true posteriors being $0.38$ bits. The actual $\mathcal{D}_{\mathrm{KL}}$ estimated with margarine is $0.05_{-0.52}^{+4.02}$.
@@
Figure 3: The graph shows how the upper limit value of $\mathcal{D}_{\mathrm{KL}}$, defined by eq. 32, changes with the standard deviation of the Gaussian random noise in the data and the RMSE error on the emulator. The dashed red line and dashed green line show the contours corresponding to the mean and 95th percentile errors for the globalemu emulator used in this work. From the intersection between these lines and the dotted vertical lines at $\sigma=5,25,50$ and 250 mK one can put an approximate upper bound on the $\mathcal{D}_{\mathrm{KL}}$ between the posterior recovered when using ARES and the emulator. These upper bounds are reported in table 3 with the bound from the 95th percentile being more conservative than from the mean RMSE across the test data. The purple scatter points show estimates of the KL divergence between the recovered posteriors for the three different noise levels. We see that even when $\sigma=5$ mK the emulated posterior is very close to the true posterior recovered with ARES and that the upper limit defined in eq. 32 while not perfect provides a good gauge on the expected KL for a given emulator error. The KL values shown in purple are also reported in table 3. In this example, the units on $\sigma$ and RMSE are given in mK, but we stress that the discussion in this paper is applicable beyond 21-cm cosmology.
@@
Table 3: The table shows the predicted upper limits on $\mathcal{D}_{\mathrm{KL}}$ for three different noise levels $\sigma=5,25,50$ and 250 mK using the mean and 95th percentile RMSE values for the emulator used in this work. The upper limits come from eq. 32 and we also report the $\mathcal{D}_{\mathrm{KL}}$ values estimated with the code margarine. We find that the actual $\mathcal{D}_{\mathrm{KL}}$ are largely consistent with the bounds and zero within error.
Since we used the same fiducial signal as DJ23 we should be able to visually compare the posteriors recovered with ARES. The comparison is complicated by the inclusion of constraints from the UV luminosity function in DJ23, which will constrain the parameters governing star formation, but we can focus on the four parameters that are largely constrained by the sky-averaged 21-cm signal namely $f_{\mathrm{esc}}$, $C_{X}$, $T_{\mathrm{min}}$ and $\log N_{\mathrm{HI}}$. There is a similarity between the posteriors for the 5 mK noise in our work and the 25 mK noise in DJ23. Similarly, consistent posteriors are seen on these four parameters when comparing our 50 mK case with the DJ23 250 mK case. DJ23 shows the posteriors for the 50 mK case without including constraints from the UVLF allowing us to compare the posteriors for all the parameters in this case. Again, we see a similarity between our 25 mK case and the DJ23 50 mK case. In the DJ23 250 mK case, the 1D posteriors for $f_{\mathrm{esc}}$ and $T_{\mathrm{min}}$ appear to be quite well constrained, and they place an upper bound on the value of $C_{X}$. We do not see as strong constraints in our 250 mK posteriors (see appendix A).
@@ 4 Discussion and Conclusions
The upper bound on the KL divergence given in eq. 32 can be used to estimate how accurate a neural network emulator needs to be for accurate posterior recovery. While we have motivated this with an example in 21-cm cosmology, emulators are being widely used in a variety of fields and the discussion outlined here is more widely applicable.
DJ23 concludes, based on the ‘emulator bias’ in eq. 37, that the recovered posteriors are inconsistent when using ARES and globalemu for data $\sigma\leq 25$ mK. However, we find that consistent constraints can be recovered using globalemu even for $\sigma=5$ mK and frame the comparison in terms of the KL divergence. A similar level of agreement was found in Breitman et al. (2024) when comparing the posteriors recovered from an analysis of the HERA 21-cm power spectrum upper limit with the emulator 21cmEMU and the semi-numerical simulation code 21cmFAST. Direct comparison with DJ23 is challenging because they included the UVLF in their analysis. The analysis and code from this paper is available at https://github.com/htjb/validating_posteriors.
The authors of DJ23 recently introduced a new emulator to the field, named 21cmLSTM. This emulator, based on Long Short-Term Memory Neural Networks (Dorigo Jones et al., 2024), represents a significant improvement in accuracy over the current state-of-the-art emulators. They do not repeat the analysis done in DJ23 but state that an ‘emulation error $<1$ mK is needed to sufficiently exploit optimistic or standard measurements of the 21-cm signal and obtain unbiased posteriors’. While similar in its heuristic nature to previous claims about required levels of accuracy (e.g. that in Bevins et al., 2021b), this claim may be influenced by the results presented in DJ23. This work provides a theoretical way to motivate the required level of emulator accuracy needed to recover an unbiased posterior estimate. The theoretical estimate supports the intuition laid out in Bevins et al. (2021b) that the error in the emulator should be $\lesssim 10\%$ of the expected noise in the data.
Although this may not have been the intention, DJ23 raised some questions about the constraints on the properties of the first galaxies and the early universe derived using globalemu from a number of works (e.g. Bevins et al., 2022c, b, 2024; Pochinda et al., 2024; Gessey-Jones et al., 2024) but also works that have used other emulators such as 21cmGEM (e.g. Monsalve et al., 2019) and power spectrum emulators (e.g. Abdurashidova et al., 2022; HERA Collaboration et al., 2023). The theoretical arguments laid out in this paper, the experimental results and the availability of the code associated with this work should reaffirm confidence in these constraints and indeed in our ability to use emulators in 21-cm cosmology and beyond.
@@ Appendix A Additional Results
In fig. 4, fig. 5 and fig. 6 we show the true and emulated posteriors recovered when $\sigma=5,50$ and 250 mK. For $\sigma=5$ mK we are able to accurately recover the true posterior when using globalemu.
Figure 4: The figure shows the recovered posteriors when modelling the data directly with ARES in blue and with globalemu in orange for $\sigma=5$ mK. As with fig. 2 there is a similarity between the two posteriors and although the upper limit on the $\mathcal{D}_{\mathrm{KL}}=9.60$ bits the calculated $\mathcal{D}_{\mathrm{KL}}$ is $0.25_{-0.25}^{+4.45}$.
Figure 5: The true posterior recovered with ARES in blue and in orange posterior recovered with the globalemu emulator for $\sigma=50$ mK. As expected from eq. 32 and fig. 3 the posterior distributions look even more alike than when the noise is 5 and 25 mK. The estimated $\mathcal{D}_{\mathrm{KL}}\leq 0.10$ based on the mean emulator RMSE or $\mathcal{D}_{\mathrm{KL}}\leq 0.97$ based on the 95th percentile emulator RMSE. The calculated $\mathcal{D}_{\mathrm{KL}}=0.09_{-0.03}^{+1.62}$.

### v2

@@ 3.5 Likelihood functions
where $N_{d}$ is the number of data points, $D$ is the mock data containing the fiducial ARES signal and Gaussian random noise with standard deviation $\sigma$, $\mathcal{M}$ is the model of the 21-cm signal from ARES or the emulated version from globalemu  [^3] The log-evidences found for the fits in this paper differ to those found in DJ23 by around 2000 log units. We do not report them here but they can be found with the chains and the code on github. We believe the difference is because we included the normalisation term on our log-likelihood, $\frac{N_{d}}{2}\log(2\pi\sigma^{2})$, and DJ23 did not. For $N_{d}=490$ and $\sigma=25$ mK then this term is equal to approximately 2027-2027. . We are assuming that the errors are not correlated across frequency, which is a common assumption in the field and was made in DJ23.
Usually, the magnitude of the noise in ones’ data is not known. Theoretical assumptions can be made about the form and magnitude of the noise using eq.˜36, however, we often fit $\sigma$ as a free parameter in the analysis. Recent works (Gessey-Jones et al., 2024; Pochinda et al., 2024) have included an additional emulator error term in their likelihood functions, such that $\sigma^{2}=\sigma_{\mathrm{instrument}}^{2}+\sigma_{\mathrm{emulator}}^{2}$. The emulator error is often estimated from the test data sets and fixed in the analysis, whereas the instrument error is fitted for. For a sky-averaged 21-cm experiment, the instrument noise is expected to be around 25 mK (e.g. Bowman et al., 2018; de Lera Acedo et al., 2022) and emulator errors are typically around 1 mK and therefore $\sigma$ is generally dominated by the contribution from the instrument.
@@ 3.6 KL divergence
We use normalising flows (NFs) implemented with margarine (Bevins et al., 2022c, 2023) to evaluate the log-probabilities $\log P$ and $\log P_{\epsilon}$ for samples $\theta\sim P(\theta|D,\mathcal{M})$. NFs are a class of generative density estimation tools that use neural networks to parameterize invertible transformations between a known distribution such as a multivariate standard normal and a more complex target distribution like $P$ and $P_{\epsilon}$. Once trained, the flows can be used to generate samples on the distributions, i.e. $\theta\sim P(\theta|D,\mathcal{M})$, and via the change of variables formula calculate log-probabilities. Errors on the KL estimates are calculated using the method detailed in Bevins et al. (2023).
For each normalising flow used in this work, we have a learning rate of $10^{-3}$ and five neural networks chained together, each with one hidden layer of 250 nodes. We train for a maximum of 1000 epochs and early stop if the validation loss has not decreased after 20 epochs, rolling back to the optimum model. For more details on the particular implementation of normalising flows used in this work, see Bevins et al. (2022c, 2023).
When estimating KL divergence from samples, as is done here, we note that it can be oversensitive to the tails of the distributions if they have been poorly sampled. Equally, if the support of $P_{\epsilon}$ does not overlap with the support of $P$ the KL divergence becomes numerically unstable and equal to infinity. We note that this does not affect the limit defined in eq.˜32 only the numerical approximations made in the following sections to illustrate the utility of the limit. As previously discussed, in general the KL divergence between $P$ and $P_{\epsilon}$ is unattainable because the true posterior is too computationally expensive.
@@ 3.7 Results
Using the trained emulator, the likelihood function with a fixed noise and margarine to calculate the KL divergence, as discussed in the previous sections, we are able to compare the posteriors recovered when using ARES and globalemu to model the signal. We use the nested sampling implementation Polychord (Handley et al., 2015b, a) to sample the likelihood function. In all the fits presented, we use the default Polychord settings [^4] See https://github.com/htjb/validating_posteriors. and we show in appendix˜C that this is sufficient to recover consistent results. As in DJ23, we test recovery when fitting a fiducial 21-cm signal with Gaussian distributed noise that has a standard deviation of 5, 25, 50 and 250 mK. We show the posteriors for 25 mK noise in fig.˜2 and for 5, 50 and 250 mK in appendix˜A.
@@
From fig.˜1, we might expect the recovery of the 1D posteriors on $f_{\mathrm{esc}}$, $C_{X}$ and $\log N_{\mathrm{HI}}$ to be worse than other parameters as the emulator error is worse at lower redshifts. However, we find that the posteriors on these parameters are very consistent, suggesting that the performance of the emulator even in this part of the parameter space is good. We report the ‘emulator bias’ for each parameter and each level of noise in appendix˜B. We find that the emulation bias is consistently below 1 for all the parameters and all the noise levels. At most, the emulation bias is 0.69 for $\log T_{\mathrm{min}}$ when $\sigma=5$ mK and we find that the average emulation bias decreases with increasing noise from 0.33 to 0.04 at 250 mK.
@@
Using eq.˜32 we can estimate an approximate upper limit on the $\mathcal{D}_{\mathrm{KL}}$ between the two posteriors for a given level of noise in the data and a given emulator error. In table˜3 we report these limits using the mean and 95th percentile RMSE values for the emulator for each level of noise considered in this paper. Since the 95th percentile RMSE is larger than the mean value and the limit is $\propto\mathrm{RMSE}^{2}$ it gives a more conservative limit on the accuracy of the emulated posterior. We also estimate the $\mathcal{D}_{\mathrm{KL}}$ between the posteriors using the method described in section˜3.6 and report these values in table˜3. We see that for all the different noise levels the $\mathcal{D}_{\mathrm{KL}}$ are approximately consistent, within uncertainty, with 0 and consistent with the upper bounds from eq.˜32.
@@
In fig.˜3 we show how the upper limit on $\mathcal{D}_{\mathrm{KL}}$ changes with the magnitude of the noise in the data and the RMSE error on the emulator as a series of blue contours. We also show contours for the mean and 95th percentile RMSE values for the emulator used in this paper as red and green dashed lines. The vertical dotted lines mark the 5, 25, 50 and 250 mK noise values and the intersection between these lines and the dashed red and green lines give the predicted upper limits on the $\mathcal{D}_{\mathrm{KL}}$ reported in table˜3. The purple scatter points give the $\mathcal{D}_{\mathrm{KL}}$ values, estimated with margarine, for each pair of posteriors. Although not perfect, it is clear that the limit in eq.˜32 provides an approximate estimate for the maximum value of the $\mathcal{D}_{\mathrm{KL}}$ we might expect for a given noise and RMSE, and hence a good guide on the accuracy required of emulators for inference. For each pair of posteriors, the estimated $\mathcal{D}_{\mathrm{KL}}$ is approximately consistent with zero.
@@
Figure 2: The posteriors recovered when fitting the fiducial ARES signal directly with ARES in blue and globalemu in orange for $\sigma=25$ mK. The lower half of the triangle plot shows a kernel density estimation (KDE) of the 2D marginalised posteriors, the diagonal shows the 1D KDEs and the upper half shows the samples. We show the fiducial parameter values as red dashed lines for reference, but stress that we are more interested in the similarity between the posteriors in this work. While there are some small difference between the posteriors, they are visually similar. However, given the mean RMSE for the emulator and the limit outlined in eq.˜32 this is not so surprising with the maximum predicted $\mathcal{D}_{\mathrm{KL}}$ between the emulated and true posteriors being $0.38$ bits. The actual $\mathcal{D}_{\mathrm{KL}}$ estimated with margarine is $0.05_{-0.52}^{+4.02}$.
@@
Figure 3: The graph shows how the upper limit value of $\mathcal{D}_{\mathrm{KL}}$, defined by eq.˜32, changes with the standard deviation of the Gaussian random noise in the data and the RMSE error on the emulator. The dashed red line and dashed green line show the contours corresponding to the mean and 95th percentile errors for the globalemu emulator used in this work. From the intersection between these lines and the dotted vertical lines at $\sigma=5,25,50$ and 250 mK one can put an approximate upper bound on the $\mathcal{D}_{\mathrm{KL}}$ between the posterior recovered when using ARES and the emulator. These upper bounds are reported in table˜3 with the bound from the 95th percentile being more conservative than from the mean RMSE across the test data. The purple scatter points show estimates of the KL divergence between the recovered posteriors for the three different noise levels. We see that even when $\sigma=5$ mK the emulated posterior is very close to the true posterior recovered with ARES and that the upper limit defined in eq.˜32 while not perfect provides a good gauge on the expected KL for a given emulator error. The KL values shown in purple are also reported in table˜3. In this example, the units on $\sigma$ and RMSE are given in mK, but we stress that the discussion in this paper is applicable beyond 21-cm cosmology.
@@
Table 3: The table shows the predicted upper limits on $\mathcal{D}_{\mathrm{KL}}$ for three different noise levels $\sigma=5,25,50$ and 250 mK using the mean and 95th percentile RMSE values for the emulator used in this work. The upper limits come from eq.˜32 and we also report the $\mathcal{D}_{\mathrm{KL}}$ values estimated with the code margarine. We find that the actual $\mathcal{D}_{\mathrm{KL}}$ are largely consistent with the bounds and zero within error.
Since we used the same fiducial signal as DJ23 we should be able to visually compare the posteriors recovered with ARES. The comparison is complicated by the inclusion of constraints from the UV luminosity function in DJ23, which will constrain the parameters governing star formation, but we can focus on the four parameters that are largely constrained by the sky-averaged 21-cm signal namely $f_{\mathrm{esc}}$, $C_{X}$, $T_{\mathrm{min}}$ and $\log N_{\mathrm{HI}}$. There is a similarity between the posteriors for the 5 mK noise in our work and the 25 mK noise in DJ23. Similarly, consistent posteriors are seen on these four parameters when comparing our 50 mK case with the DJ23 250 mK case. DJ23 shows the posteriors for the 50 mK case without including constraints from the UVLF allowing us to compare the posteriors for all the parameters in this case. Again, we see a similarity between our 25 mK case and the DJ23 50 mK case. In the DJ23 250 mK case, the 1D posteriors for $f_{\mathrm{esc}}$ and $T_{\mathrm{min}}$ appear to be quite well constrained, and they place an upper bound on the value of $C_{X}$. We do not see as strong constraints in our 250 mK posteriors (see appendix˜A).
@@ 4 Discussion and Conclusions
The upper bound on the KL divergence given in eq.˜32 can be used to estimate how accurate a neural network emulator needs to be for accurate posterior recovery. While we have motivated this with an example in 21-cm cosmology, emulators are being widely used in a variety of fields and the discussion outlined here is more widely applicable.
DJ23 concludes, based on the ‘emulator bias’ in eq.˜37, that the recovered posteriors are inconsistent when using ARES and globalemu for data $\sigma\leq 25$ mK. However, we find that consistent constraints can be recovered using globalemu even for $\sigma=5$ mK and frame the comparison in terms of the KL divergence. A similar level of agreement was found in Breitman et al. (2024) when comparing the posteriors recovered from an analysis of the HERA 21-cm power spectrum upper limit with the emulator 21cmEMU and the semi-numerical simulation code 21cmFAST. Direct comparison with DJ23 is challenging because they included the UVLF in their analysis. The analysis and code from this paper is available at https://github.com/htjb/validating_posteriors.
The authors of DJ23 recently introduced a new emulator to the field, named 21cmLSTM. This emulator, based on Long Short-Term Memory Neural Networks (Dorigo Jones et al., 2024), represents a significant improvement in accuracy over the current state-of-the-art emulators. They do not repeat the analysis done in DJ23 but state that an ‘emulation error <1<1 mK is needed to sufficiently exploit optimistic or standard measurements of the 21-cm signal and obtain unbiased posteriors’. While similar in its heuristic nature to previous claims about required levels of accuracy (e.g. that in Bevins et al., 2021b), this claim may be influenced by the results presented in DJ23. This work provides a theoretical way to motivate the required level of emulator accuracy needed to recover an unbiased posterior estimate. The theoretical estimate supports the intuition laid out in Bevins et al. (2021b) that the error in the emulator should be $\lesssim 10\%$ of the expected noise in the data.
Although this may not have been the intention, DJ23 raised some questions about the constraints on the properties of the first galaxies and the early universe derived using globalemu from a number of works (e.g. Bevins et al., 2022a, b, 2024; Pochinda et al., 2024; Gessey-Jones et al., 2024) but also works that have used other emulators such as 21cmGEM (e.g. Monsalve et al., 2019) and power spectrum emulators (e.g. Abdurashidova et al., 2022; HERA Collaboration et al., 2023). The theoretical arguments laid out in this paper, the experimental results and the availability of the code associated with this work should reaffirm confidence in these constraints and indeed in our ability to use emulators in 21-cm cosmology and beyond.
@@ Appendix A Additional Results
In fig.˜4, fig.˜5 and fig.˜6 we show the true and emulated posteriors recovered when $\sigma=5,50$ and 250 mK. For $\sigma=5$ mK we are able to accurately recover the true posterior when using globalemu.
Figure 4: The figure shows the recovered posteriors when modelling the data directly with ARES in blue and with globalemu in orange for $\sigma=5$ mK. As with fig.˜2 there is a similarity between the two posteriors and although the upper limit on the $\mathcal{D}_{\mathrm{KL}}=9.60$ bits the calculated $\mathcal{D}_{\mathrm{KL}}$ is $0.25_{-0.25}^{+4.45}$.
Figure 5: The true posterior recovered with ARES in blue and in orange posterior recovered with the globalemu emulator for $\sigma=50$ mK. As expected from eq.˜32 and fig.˜3 the posterior distributions look even more alike than when the noise is 5 and 25 mK. The estimated $\mathcal{D}_{\mathrm{KL}}\leq 0.10$ based on the mean emulator RMSE or $\mathcal{D}_{\mathrm{KL}}\leq 0.97$ based on the 95th percentile emulator RMSE. The calculated $\mathcal{D}_{\mathrm{KL}}=0.09_{-0.03}^{+1.62}$.